### 07 Jan 2024 v5

@@ Abstract
Recently, learned image compression has achieved impressive performance. The entropy model, which estimates the distribution of the latent representation, plays a crucial role in enhancing rate-distortion performance. The latent representation contains channel-wise, local spatial, and global spatial correlations. However, existing global context modules rely on computationally intensive quadratic complexity computations to capture global correlations. The quadratic complexity imposes limitations on the potential of high-resolution image coding. Moreover, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To address these limitations, we propose the Linear Complexity Attention-based Multi-Reference Entropy Model (${}^{++}$). ${}^{++}$ effectively captures the diverse range of correlations inherent in the latent representation. Specifically, the latent representation is first divided into multiple slices. When compressing a particular slice, the previously compressed slices serve as its channel-wise contexts. To capture local contexts without sacrificing performance, we introduce a novel checkerboard attention module with linear complexity. Additionally, to capture global contexts, we propose the linear complexity attention-based global correlations capturing by leveraging the decomposition of the softmax operation. The attention map of the previously decoded slice is implicitly computed and employed to predict global correlations in the current slice. Based on ${}^{++}$, we propose image compression model ${}^{++}$. Extensive experimental evaluations demonstrate that our ${}^{++}$ achieves state-of-the-art performance, reducing BD-rate by $13.39\%$ on the Kodak dataset compared to VTM-17.0 in Peak Signal-to-Noise Ratio (PSNR). Furthermore, ${}^{++}$ exhibits linear GPU memory consumption with resolution, making it highly suitable for high-resolution image coding. Code and pre-trained models are available at https://github.com/JiangWeibeta/MLIC.
Index Terms:
Entropy model, Learned image compression.
I Introduction
Due to the rise of social media, tens of millions of images are generated and transmitted on the web every second. In order to conserve bandwidth, service providers are compelled to seek more efficient and effective image compression methods. Although traditional coding methods like JPEG [1], JPEG2000 [2], AVC [3], HEVC [4], and VVC [5] have achieved commendable performance, their design relies on manual design for each module. This lack of joint optimization hampers their ability to fully exploit the potential for further advancements in image compression.
Recently, various learned image compression models  [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] have emerged, showcasing impressive performance gains.
@@ I Introduction
Figure 1: BD-Rate-GPU Memory Consumption during inference on CLIC Professional Valid [22] with 2K resolution. Our ${}^{++}$ achieves a better trade-off between performance and GPU memory consumption.
@@
Notably, certain learned image compression models [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 27, 36] are already comparable to the advanced traditional method VVC. These models predominantly rely on auto-encoders or variational auto-encoders [37], and follow a process that involves transform, quantization, entropy coding, and inverse transform. Entropy coding plays an important role in boosting model performance. An entropy model is utilized to estimate the entropy of the latent representation. A powerful and accurate entropy model usually leads to fewer bits. Expanding contexts of entropy model in learned codecs plays the same role as expanding prediction modes in traditional codecs.
@@
State-of-the-art learned image compression models [24, 23, 28, 38, 29, 30, 36] commonly enhance the entropy model by incorporating a hyper-prior module [17] or a context module [19]. These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consumes huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity attention-based multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as ${}^{++}$ to differ from our prior work [30] presented at ACMMM 2023. Based on ${}^{++}$, we introduce ${}^{++}$, which achieve state-of-the-art performance.
@@
In our approach, the latent representations is divided into multiple slices along the channel dimension. When compressing a particular slice, the previously compressed slices serve as its channel-wise contexts, which are extract by a channel-wise context module. Local and global context modeling are conducted separately for each slice. The utilization of an auto-regressive local context module [19, 39] leads to serial decoding, while a checkerboard context module [40] facilitates two-pass parallel decoding by dividing the latent representations into anchor and non-anchor parts. However, it is worth noting that the checkerboard context module may result in a performance degradation of up to $4\%$ [41]. To address this issue, we propose a novel overlapped checkerboard window attention with linear complexity, which further enhance the local context capturing while retaining two-pass decoding. Some previous methods focus on global context modeling [26, 14], which typically involve quadratic complexity or the utilization of additional bits to store global similarity as side information. Additionally, these global context modules often collaborate with serial local spatial context modules, further increasing the computational complexity. Assuming comparable spatial correlations across different slices, we initially calculate the attention map of the previously decoded $i-1$-th slice in a vanilla approach. This attention map is utilized to predict the global correlations within the $i$-th slice. However, in the vanilla attention mechanism, the softmax operation dictates the order of computation among tensors, where the attention map, product of queries and keys are required to be computed first. To circumvent the quadratic complexity, we employ the decomposition of a softmax operation into two independent softmax operations such that the product of keys and values can be computed first, resulting in linear complexity. The proposed linear complexity attention-based global context modules capture global correlations in an implicit way, as there is no need to directly compute the attention map during training and testing. In addition, we also propose the linear complexity inter-slice global spatial context modules to explore the global correlations in all preceding slices. The linear complexity context allows our model to have a linear relationship between consumed GPU memory and resolution without additional bits, while having the performance gain that comes from the global contexts. Ultimately, we integrate the channel, local, intra-slice, inter-slice global contexts, along with the side information for multi-reference entropy modeling. Our contributions are summarized as follows:
@@
  • We design linear complexity attention-based multi-reference entropy model ${}^{++}$ which captures local spatial, global spatial and channel contexts, as well as hyper-prior side information. Based on ${}^{++}$, we propose ${}^{++}$, which achieves state-of-the-art performance. The complexity and GPU memory consumption of our model is linear with the resolution. Our proposed ${}^{++}$ achieved a better trade-off between complexity and performance as depicted in Fig. 1 .
In comparison to our previous work presented at ACMMM 2023 [30], ${}^{++}$ introduces several significant advancements. The primary distinction lies in the utilization of the proposed linear complexity global spatial context modules without sacrificing performance, as opposed to the quadratic complexity observed in our prior work. This achievement is primarily attributed to the division of the softmax operation, which eliminates the need for a specific order of tensor computation. Our proposed modules incorporate advanced techniques such as learnable position embedding and depth-wise residual bottlenecks [36]. Furthermore, ${}^{++}$ captures inter-slice global correlations from all previous slices, in contrast to our previous work [30], which only considers correlations within the previous one slice. ${}^{++}$ exhibits several advantages, including reduced GPU memory consumption and faster encoding and decoding speed. The linear complexity makes our ${}^{++}$ highly suitable for high-resolution image coding. These advancements in ${}^{++}$ contribute to the field of image compression by offering improved efficiency and performance, while maintaining high-quality compression capabilities.
II Related Works
II-A Learned Image Compression
Learned image compression [15, 16, 7] aims to optimize the trade-off between distortion $\mathcal{D}$ and entropy, where entropy is typically measured in terms of bit-rate $\mathcal{R}$. Large bit-rate usually leads to lower distortion . Lagrange multiplier $\lambda$ is employed to adjust the weight of distortion to control the target bit-rate. The optimization target is
@@ II-A Learned Image Compression
The fundamental learned image compression framework [15, 16] is based on the an auto-encoder with a rate penalty. This framework comprises an analysis transform $g_{a}$, a quantization function $Q$, a synthesis transform $g_{s}$ and an entropy model to estimate rates. The process can be formulated as:
$\bm{y}=g_{a}(\bm{x};\theta),\hat{\bm{y}}=Q(\bm{y}),\hat{\bm{x}}=g_{s}(\hat{\bm{y}};\phi),$ (2)
@@
where $\bm{x}$ represents the input image, $g_{a}$ transform the $\bm{x}$ to compact latent representation $\bm{y}$. $\bm{y}$ is quantized to $\hat{\bm{y}}$ for entropy coding. $\hat{\bm{x}}$ represents the decompressed image. $\theta$ and $\phi$ are parameters of $g_{a}$ and $g_{s}$. Since quantization is non-differentiable, it can be addressed during training by either adding uniform noise $\mathcal{U}(-0.5,0.5)$ [15, 17] or using the straight-through estimator (STE) [16]. In particular, when uniform noise is added, the rate-distortion optimization target in Equation 1 is equivalent to evidence lower bound (ELBO) optimization in variational auto-encoders [37]. To enhance non-linearity, Generalized Divisive Normalization (GDN) [42] layers or its variants [26, 8] are employed. Additionally, self-attention [43, 44, 45, 25, 27, 10, 14], ensemble techniques [9], and block partition [6] are utilized in transform modules for more compact latent representations. In the basic model, a factorized or a non-adaptive density entropy model is adopted.
@@
In subsequent works, a hyper-prior module [17] is introduced to extract side information $\hat{\bm{z}}$ from $\bm{y}$. The hyper-prior model estimates the distribution of $\hat{\bm{y}}$ from $\hat{\bm{z}}$. A univariate Gaussian distribution is commonly employed for the hyper-prior. Some works extend it to a mean and scale Gaussian distribution [19], asymmetric Gaussian distribution [46], Gaussian mixture model [24, 47], and Gaussian-Laplacian-Logistic mixture model [48] for more flexible distribution modeling.
II-B Context-based Entropy Modeling
@@
Numerous approaches [19, 23, 26] have been proposed to improve the accuracy of context modeling in learned image compression. These methods encompass various types of context modules, including local spatial, global spatial, and channel-wise context modules.
@@
Local spatial context modules aim to capture correlations between adjacent symbols. For instance, Minnen et al. [19] utilize a pixel-cnn-like [39] masked convolutional layer to capture local correlations between $\hat{\bm{y}}_{i}$ and symbols $\hat{\bm{y}}_{<i}$, resulting in serial decoding. He et al. [40] divide latent representation $\hat{\bm{y}}$ into anchor part $\hat{\bm{y}}_{a}$ and non-anchor part $\hat{\bm{y}}_{na}$, employing a checkerboard convolution to extract contexts of $\hat{\bm{y}}_{na}$ from $\hat{\bm{y}}_{a}$, thereby achieving two-pass parallel decoding.
@@
On the other hand, some approaches focus on modeling correlations between distant symbols. In [26], neighboring left and top symbols serve as bases for computing the similarity between the target symbol and its previous symbols. Guo et al. [14] employ the $L2$ distances of symbols to predict global casual dependencies among symbols. In [49], the side information is divided into global side information and local side information, introducing additional bits. However, these global context modules are typically combined with serial auto-regressive context modules, which further increase decoding latency. Moreover, existing global context modules [26, 14, 30] often exhibit quadratic complexity, making them challenging to apply in high-resolution image coding. Alternatively, they rely on extra side information [49], which increases the bit-rate.
@@
Minnen et al. [23] model contexts between channels. $\hat{\bm{y}}$ is evenly divided to slices. The current slice $\hat{\bm{y}}^{i}$ is conditioned on previously decoded slices $\hat{\bm{y}}^{<i}$. To address the uneven distribution of information among slices, an unevenly grouped channel-wise context module is introduced in [29].
@@
While some local and channel-wise context modules [50, 29, 30] have demonstrated impressive performance, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. Addressing these correlations has the potential to further enhance the performance of image compression models.
@@
Figure 2: Visualization of channels of latent representation of Kodim19 extracted by Cheng’20 [24] (optimized for MSE, $\lambda=0.0483$) to illustrate channel-wise redundancy.
@@
Figure 3: Heatmap of spatial cosine similarity of latent representation of Kodim19 extracted by Cheng’20 [24] (optimized for MSE, $\lambda=0.0483$) to visualize global spatial and local spatial redundancy.
@@
$\bm{x},\hat{\bm{x}}$ Input and decoded image
$\bm{y},\hat{\bm{y}}$ Non-quantized and quantized latent representation
$\hat{\bm{y}}^{i}$ The $i$-th slice of $\hat{\bm{y}}$
$\hat{\bm{y}}_{ac},\hat{\bm{y}}_{na}$ Anchor and non-anchor part of $\hat{\bm{y}}$
$\bm{z},\hat{\bm{z}}$ Non-quantized and quantized side information
@@
$\mu,\sigma$ Mean and scale of $\hat{\bm{y}}$
@@
${\bm{\Phi}}_{h},\bm{\Phi}_{ch},{\bm{\Phi}}_{lc}$ Hyper-prior, channel-wise and local spatial context
${\bm{\Phi}}_{gc,intra}$ Intra-slice global spatial context
${\bm{\Phi}}_{gc,inter}$ Inter-slice global spatial context
MEM ${}^{++}$ Linear complexity attention-based
@@
$M,N,S$ Channel number of $\bm{y}$, $\bm{z}$, and $\hat{\bm{y}}^{i}$
@@
TABLE I: Explanations of notations.
Figure 4: The overall architecture of ${}^{++}$. $\downarrow$ means down-sampling. $\uparrow$ means up-sampling. / means stride equals $1$. Red line is the dataflow during decoding. ${\bm{x}}$ is the input image and $\hat{\bm{x}}$ is the reconstructed image. $\bm{y}$ is the latent representation and $\hat{\bm{y}}$ is the quantized latent representation. $\hat{\bm{y}}^{i}$ is the $i$-th slice of $\hat{\bm{y}}$. $\hat{\bm{y}}^{i}_{ac}$ is the anchor part of $\hat{\bm{y}}^{i}$. $\hat{\bm{y}}^{i}_{na}$ is the non-anchor part of $\hat{\bm{y}}^{i}$.
III Method
III-A Motivation

### 20 Feb 2024 v9

@@ Abstract
Recently, learned image compression has achieved impressive performance. The entropy model, which estimates the distribution of the latent representation, plays a crucial role in enhancing rate-distortion performance. The latent representation contains channel-wise, local spatial, and global spatial correlations. However, existing global context modules rely on computationally intensive quadratic complexity computations to capture global correlations. The quadratic complexity imposes limitations on the potential of high-resolution image coding. Moreover, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To address these limitations, we propose the Linear Complexity Multi-Reference Entropy Model (${}^{++}$). ${}^{++}$ effectively captures the diverse range of correlations inherent in the latent representation. Specifically, the latent representation is first divided into multiple slices. When compressing a particular slice, the previously compressed slices serve as its channel-wise contexts. To capture local contexts without sacrificing performance, we introduce a novel checkerboard attention module with linear complexity. Additionally, to capture global contexts, we propose the linear complexity attention-based global correlations capturing by leveraging the decomposition of the softmax operation. The attention map of the previously decoded slice is implicitly computed and employed to predict global correlations in the current slice. Based on ${}^{++}$, we propose image compression model ${}^{++}$. Extensive experimental evaluations demonstrate that our ${}^{++}$ achieves state-of-the-art performance, reducing BD-rate by $13.39\%$ on the Kodak dataset compared to VTM-17.0 in Peak Signal-to-Noise Ratio (PSNR). Furthermore, ${}^{++}$ exhibits linear GPU memory consumption with resolution, making it highly suitable for high-resolution image coding. Code and pre-trained models are available at https://github.com/JiangWeibeta/MLIC.
Machine Learning, ICML
wei.jiang1999@outlook.com, jiangwei@stu.pku.edu.cn
rgwang@pkusz.edu.cn
1 Introduction
Due to the rise of social media, tens of millions of images are generated and transmitted on the web every second.
@@ 1 Introduction
Figure 1: BD-Rate-GPU Memory Consumption during inference on CLIC Professional Valid (Toderici et al., 2020) with 2K resolution. Our ${}^{++}$ achieves a better trade-off between performance and GPU memory consumption.
In order to conserve bandwidth, service providers are compelled to seek more efficient and effective image compression methods. Although traditional coding methods like JPEG (Pennebaker & Mitchell, 1992), JPEG2000 (Charrier et al., 1999), AVC (Wiegand et al., 2003), HEVC (Sullivan et al., 2012), and VVC (Bross et al., 2021) have achieved commendable performance, their design relies on manual design for each module. This lack of joint optimization hampers their ability to fully exploit the potential for further advancements in image compression.
@@
Recently, various learned image compression models  (Wu et al., 2022; Guo et al., 2022; Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Hu et al., 2020; Ma et al., 2020) have emerged, showcasing impressive performance gains. Notably, certain learned image compression models (Minnen & Singh, 2020; Cheng et al., 2020; Zou et al., 2022; Qian et al., 2020; Xie et al., 2021; He et al., 2022; Jiang et al., 2023b; Chen et al., 2021; Gao et al., 2021; Chen et al., 2022; Koyuncu et al., 2022; Duan et al., 2023; Liu et al., 2023; Jiang et al., 2023a; Fu et al., 2023a; Tang et al., 2023; Chen & Ma, 2023) are already comparable to the advanced traditional method VVC. These models predominantly rely on auto-encoders or variational auto-encoders (Kingma & Welling, 2014), and follow a process that involves transform, quantization, entropy coding, and inverse transform. Entropy coding plays an important role in boosting model performance. An entropy model is utilized to estimate the entropy of the latent representation. A powerful and accurate entropy model usually leads to fewer bits. Expanding contexts of entropy model in learned codecs plays the same role as expanding prediction modes in traditional codecs.
@@
State-of-the-art learned image compression models (Cheng et al., 2020; Minnen & Singh, 2020; Xie et al., 2021; Zhu et al., 2022b; He et al., 2022; Jiang et al., 2023b, a) commonly enhance the entropy model by incorporating a hyper-prior module (Ballé et al., 2018) or a context module (Minnen et al., 2018). These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consume huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as ${}^{++}$ to differ from our prior work (Jiang et al., 2023b) presented at ACMMM 2023. Based on ${}^{++}$, we introduce ${}^{++}$, which achieve state-of-the-art performance.
@@
In our approach, the latent representations is divided into multiple slices along the channel dimension. When compressing a particular slice, the previously compressed slices serve as its channel-wise contexts, which are extracted by a channel-wise context module. Local and global context modeling are conducted separately for each slice. The utilization of an auto-regressive local context module (Minnen et al., 2018; Van den Oord et al., 2016) leads to serial decoding, while a checkerboard context module (He et al., 2021) facilitates two-pass parallel decoding by dividing the latent representations into anchor and non-anchor parts. However, it is worth noting that the checkerboard context module may result in a performance degradation of up to $4\%$ (Qian et al., 2022). To address this issue, we propose a novel overlapped checkerboard window attention with linear complexity, which further enhances the local context capturing while retaining two-pass decoding. Some previous methods focus on global context modeling (Qian et al., 2020; Guo et al., 2022), which typically involve quadratic complexity or the utilization of additional bits to store global similarity as side information. Additionally, these global context modules often collaborate with serial local spatial context modules, further increasing the computational complexity. Assuming comparable spatial correlations across different slices, we initially calculate the attention map of the previously decoded $i-1$-th slice in a vanilla approach. This attention map is utilized to predict the global correlations within the $i$-th slice. However, in the vanilla attention mechanism, the softmax operation dictates the order of computation among tensors, where the attention map, product of queries and keys are required to be computed first. To circumvent the quadratic complexity, we employ the decomposition of a softmax operation into two independent softmax operations such that the product of keys and values can be computed first, resulting in linear complexity. The proposed linear complexity attention-based global context modules capture global correlations in an implicit way, as there is no need to directly compute the attention map during training and testing. In addition, we also propose the linear complexity inter-slice global spatial context modules to explore the global correlations in all preceding slices. The linear complexity context allows our model to have a linear relationship between consumed GPU memory and resolution without additional bits, while having the performance gain that comes from the global contexts. Ultimately, we integrate the channel, local, intra-slice, inter-slice global contexts, along with the side information for multi-reference entropy modeling. Our contributions are summarized as follows:
@@
  • We design linear complexity multi-reference entropy model ${}^{++}$ which captures local spatial, global spatial and channel contexts, as well as hyper-prior side information. Based on ${}^{++}$, we propose ${}^{++}$, which achieves state-of-the-art performance. The complexity and GPU memory consumption of our model is linear with the resolution. Our proposed ${}^{++}$ achieved a better trade-off between complexity and performance as depicted in Fig. 1 .
In comparison to our previous work presented at ACMMM 2023 (Jiang et al., 2023b), ${}^{++}$ introduces several significant advancements. The primary distinction lies in the utilization of the proposed linear complexity global spatial context modules without sacrificing performance, as opposed to the quadratic complexity observed in our prior work. This achievement is primarily attributed to the division of the softmax operation, which eliminates the need for a specific order of tensor computation. Our proposed modules incorporate advanced techniques such as learnable position embedding and depth-wise residual bottlenecks (Jiang et al., 2023a). Furthermore, ${}^{++}$ captures inter-slice global correlations from all previous slices, in contrast to our previous work (Jiang et al., 2023b), which only considers correlations within the previous one slice. ${}^{++}$ exhibits several advantages, including reduced GPU memory consumption and faster encoding and decoding speed. The linear complexity makes our ${}^{++}$ highly suitable for high-resolution image coding. These advancements in ${}^{++}$ contribute to the field of image compression by offering improved efficiency and performance, while maintaining high-quality compression capabilities.
2 Related Works
2.1 Learned Image Compression
Learned image compression (Ballé et al., 2017; Theis et al., 2017) aims to optimize the trade-off between distortion $\mathcal{D}$ and entropy, where entropy is typically measured in terms of bit-rate $\mathcal{R}$. Large bit-rate usually leads to lower distortion . Lagrange multiplier $\lambda$ is employed to adjust the weight of distortion to control the target bit-rate. The optimization target is
@@ 2.1 Learned Image Compression
The fundamental learned image compression framework (Ballé et al., 2017; Theis et al., 2017) is based on the an auto-encoder with a rate penalty. This framework comprises an analysis transform $g_{a}$, a quantization function $Q$, a synthesis transform $g_{s}$ and an entropy model to estimate rates. The process can be formulated as:
$\boldsymbol{y}=g_{a}(\boldsymbol{x};\theta),\hat{\boldsymbol{y}}=Q(\boldsymbol{y}),\hat{\boldsymbol{x}}=g_{s}(\hat{\boldsymbol{y}};\phi),$ (2)
@@
where $\boldsymbol{x}$ represents the input image, $g_{a}$ transform the $\boldsymbol{x}$ to compact latent representation $\boldsymbol{y}$. $\boldsymbol{y}$ is quantized to $\hat{\boldsymbol{y}}$ for entropy coding. $\hat{\boldsymbol{x}}$ represents the decompressed image. $\theta$ and $\phi$ are parameters of $g_{a}$ and $g_{s}$. Since quantization is non-differentiable, it can be addressed during training by either adding uniform noise $\mathcal{U}(-0.5,0.5)$ (Ballé et al., 2017, 2018) or using the straight-through estimator (STE) (Theis et al., 2017). In particular, when uniform noise is added, the rate-distortion optimization target in Equation 1 is equivalent to evidence lower bound (ELBO) optimization in variational auto-encoders (Kingma & Welling, 2014). To enhance non-linearity, Generalized Divisive Normalization (GDN) (Ballé et al., 2015) layers or its variants (Qian et al., 2020) are employed. Additionally, self-attention (Vaswani et al., 2017; Liu et al., 2021; Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Guo et al., 2022), ensemble techniques (Wang et al., 2021), and block partition (Wu et al., 2022) are utilized in transform modules for more compact latent representations. In the basic model, a factorized or a non-adaptive density entropy model is adopted.
@@
In subsequent works, a hyper-prior module (Ballé et al., 2018) is introduced to extract side information $\hat{\boldsymbol{z}}$ from $\boldsymbol{y}$. The hyper-prior model estimates the distribution of $\hat{\boldsymbol{y}}$ from $\hat{\boldsymbol{z}}$. A univariate Gaussian distribution is commonly employed for the hyper-prior. Some works extend it to a mean and scale Gaussian distribution (Minnen et al., 2018), asymmetric Gaussian distribution (Cui et al., 2021), Gaussian mixture model (Cheng et al., 2020; Liu et al., 2020b), and Gaussian-Laplacian-Logistic mixture model (Fu et al., 2023b) for more flexible distribution modeling.
2.2 Context-based Entropy Modeling
@@
Numerous approaches (Minnen et al., 2018; Minnen & Singh, 2020; Qian et al., 2020) have been proposed to improve the accuracy of context modeling in learned image compression. These methods encompass various types of context modules, including local spatial, global spatial, and channel-wise context modules.
@@
Local spatial context modules aim to capture correlations between adjacent symbols. For instance, Minnen et al. (Minnen et al., 2018) utilize a pixel-cnn-like (Van den Oord et al., 2016) masked convolutional layer to capture local correlations between $\hat{\boldsymbol{y}}_{i}$ and symbols $\hat{\boldsymbol{y}}_{<i}$, resulting in serial decoding. He et al. (He et al., 2021) divide latent representation $\hat{\boldsymbol{y}}$ into anchor part $\hat{\boldsymbol{y}}_{a}$ and non-anchor part $\hat{\boldsymbol{y}}_{na}$, employing a checkerboard convolution to extract contexts of $\hat{\boldsymbol{y}}_{na}$ from $\hat{\boldsymbol{y}}_{a}$, thereby achieving two-pass parallel decoding.
@@
On the other hand, some approaches focus on modeling correlations between distant symbols. In (Qian et al., 2020), neighboring left and top symbols serve as bases for computing the similarity between the target symbol and its previous symbols. Guo et al. (Guo et al., 2022) employ the $L2$ distances of symbols to predict global casual dependencies among symbols. In (Kim et al., 2022), the side information is divided into global side information and local side information, introducing additional bits. However, these global context modules are typically combined with serial auto-regressive context modules, which further increase decoding latency. Moreover, existing global context modules (Qian et al., 2020; Guo et al., 2022; Jiang et al., 2023b) often exhibit quadratic complexity, making them challenging to apply in high-resolution image coding. Alternatively, they rely on extra side information (Kim et al., 2022), which increases the bit-rate.
@@
Minnen et al. (Minnen & Singh, 2020) model contexts between channels. $\hat{\boldsymbol{y}}$ is evenly divided to slices. The current slice $\hat{\boldsymbol{y}}^{i}$ is conditioned on previously decoded slices $\hat{\boldsymbol{y}}^{<i}$. To address the uneven distribution of information among slices, an unevenly grouped channel-wise context module is introduced in (He et al., 2022).
@@
While some local and channel-wise context modules (Ma et al., 2021; He et al., 2022; Jiang et al., 2023b) have demonstrated impressive performance, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. Addressing these correlations has the potential to further enhance the performance of image compression models.
@@
Figure 2: Visualization of channels of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to illustrate channel-wise redundancy.
@@
Figure 3: Heatmap of spatial cosine similarity of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to visualize global spatial and local spatial redundancy.
@@
$\boldsymbol{x},\hat{\boldsymbol{x}}$ Input and decoded image
$\boldsymbol{y},\hat{\boldsymbol{y}}$ Non-quantized and quantized latent representation
$\hat{\boldsymbol{y}}^{i}$ The $i$-th slice of $\hat{\boldsymbol{y}}$
$\hat{\boldsymbol{y}}_{ac},\hat{\boldsymbol{y}}_{na}$ Anchor and non-anchor part of $\hat{\boldsymbol{y}}$
$\boldsymbol{z},\hat{\boldsymbol{z}}$ Non-quantized and quantized side information
@@
$\mu,\sigma$ Mean and scale of $\hat{\boldsymbol{y}}$
@@
${\boldsymbol{\Phi}}_{h},\boldsymbol{\Phi}_{ch},{\boldsymbol{\Phi}}_{lc}$ Hyper-prior, channel-wise and local spatial context
${\boldsymbol{\Phi}}_{gc,intra}$ Intra-slice global spatial context
${\boldsymbol{\Phi}}_{gc,inter}$ Inter-slice global spatial context
MEM ${}^{++}$ Linear complexity
@@
$M,N,S$ Channel number of $\boldsymbol{y}$, $\boldsymbol{z}$, and $\hat{\boldsymbol{y}}^{i}$
@@
Table 1: Explanations of notations.
Figure 4: The overall architecture of ${}^{++}$. $\downarrow$ means down-sampling. $\uparrow$ means up-sampling. / means stride equals $1$. Red line is the dataflow during decoding. ${\boldsymbol{x}}$ is the input image and $\hat{\boldsymbol{x}}$ is the reconstructed image. $\boldsymbol{y}$ is the latent representation and $\hat{\boldsymbol{y}}$ is the quantized latent representation. $\hat{\boldsymbol{y}}^{i}$ is the $i$-th slice of $\hat{\boldsymbol{y}}$. .
3 Method
3.1 Motivation